the sides of a scalene triangle are representing by x,x-1,and x +1.The perimeter of the triangle is 18 inches. What are the lengths of lengths of the 3 sides in inches?

Answer by stanbon(75887)

the sides of a scalene triangle are representing by x,x-1,and x +1.The perimeter of the triangle is 18 inches. What are the lengths of lengths of the 3 sides in inches?
----
Equation:
x + x-1 + x+1 = 18
3x = 18
x = 6
---
side 1: x = 6
side 2: x-1 = 5
side 3: x+1 = 7
========
Cheers,
Stan H.
========